Fermi National Laboratory


Accelerator Update

Friday January 9
The day shift began with the Tevatron (Tev) in quench recovery, stacking in the Antiproton Source (Pbar), and with beam to MiniBoone.

By 11:12 AM the Tev quench recovery is complete, Tev will begin orbit smoothing studies before going into shot setup.

Tev begins shot setup at 4:25 PM.

At 6:47 PM the Tev quenched (D1, A1, and F4) just after it reached flat top.

At 7:03 PM Pbar resumes stacking.

Tev begins ramping at 11:36 PM.

Saturday January 10
The Midnight shift began with Tev studies, stacking, and with beam to MiniBoone.

At 1:29 AM the Tev quenches once again.

Quench may have been caused by a missing + 5 volts on a I/O crate at D4. The crate power supply has been changed.

The Tev is ramping once again at 4:33 AM.

At 9:33 AM the Tev begins shot setup.

At 9:42 AM it was determined the Pbar Lithium Lens is failing and will need to be replaced. The replacement will occur after the Tev store is in.

At 11:58 AM the Tev has quenched again. The quench occurred at E48.

Replacement of Lithium lens begins at 11:58 AM.

At 2:11 PM the Tev is ramping once again, studies have started to determine the causes for the quench.

Studies began in the 400 Mev line at 2:20 PM and last until 7:05 PM, no beam to Miniboone during this time.

Beam returns to MiniBoone at 7:05 PM.

At 11:40 PM the lithium lens change out is complete, Pbar has returned to stacking.

Sunday January 11
The Midnight shift began with Tev studies, stacking, and with beam to MiniBoone. At 9:23 AM the Tev begins shot setup.

At 11:34 AM, Operations established store #3148 with an initial luminosity of 36.08 E30.

Pbar returns to stacking at 11:38 AM.

At 4:24 PM it was determined that the lithium lens that was installed yesterday has failed change out of the lens has begun.

Studies began in the 400 Mev line at 8:04 PM and last until 1:42 AM, no beam to Miniboone during this time.

Monday January 12
By 3:30 AM the lithium lens has been replaced and Pbar has returned to stacking.

Today's Plans
Terminate the store at 8:30 AM, followed by Tev studies while Pbar continues to stack. Shot setup will begin sometime after 6:00 PM.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/12/2004   email Fermilab